OpenStudy (anonymous):

A spinner is divided in 8 equal sections. 5 sections are red and three are green. If the spinner spun three times what is the probability that it land on red exactly twice?

OpenStudy (anonymous):

probability it lands red on any one spin is \(\frac{3}{8}\) doesn't land red with probability \(\frac{5}{8}\) one want exactly two out of three, so it is \[3\times (\frac{3}{8})^2\times \frac{5}{8}\] meaning 2 red, one not red and the 3 because there are three ways to choose the 2 red out of 3 total

OpenStudy (anonymous):

no no no that is wrong!

OpenStudy (anonymous):

red is with probability \(\frac{5}{8}\) not red with probability \(\frac{3}{8}\) so \[3\left(\frac{5}{8}\right)^2\left(\frac{3}{8}\right)\]

OpenStudy (anonymous):

ignore first answer i read it incorrectly

OpenStudy (anonymous):

oh ok thanks a lot ... Satellite

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Yes that is right, you use the formula n C r p^r q^n-r with p as the probability the event will occur and q being the probability it wont occur.